OpenStudy (anonymous):

Simplify the following expression, and rewrite it in an equivalent form with positive exponents. -30x^4y^-4 / 120x^2y^-4 This is in the form of a fraction. Thank you for your help.

OpenStudy (bahrom7893):

-30x^4y^-4 / 120x^2y^-4 = -30x^4/120x^2 = -x^4/4x^2 = -x^2/4

OpenStudy (anonymous):

So the ys canceled out. ok. Thank you for your help. Can I put another question under this one like a thread?

OpenStudy (bahrom7893):

yea sure..

OpenStudy (anonymous):

(-33x^-2y^2 / 165x^-1)^3 I am going to work it out and could you let me know if it is right or where I messed up? Short answer: Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. =-33^3(x^-2)^3(y^2)^3 / 165^3(x^-1)^3 =-35937x^-6y^6 / 4492125x^-3 =-1/125*x^(-6- -3)*y^6 =-1/125x^-3y^6

OpenStudy (bahrom7893):

that's right!

OpenStudy (anonymous):

Wow, ok Thank you ok I have more I need help on. I will type it up, give me a minute.

OpenStudy (anonymous):

Short answer: Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. (-33xy^-1 / 66x^0y^-3)^2 This is my work: =-33^2x^2(y^-1)^2 / 66^2(x^0=1=1^2)(y^-3)^2 =1089x^2y^(-1-2) / 4356*1y^(-3-2) =1089x^2y^-3 / 4356y^-5 I do not know how to make them positive, can you help me?